I am having trouble figuring this out. How do I break down the calculations?

INSTRUCTIONS FOR MILESTONE 2 (Due Week 4) IMPORTANT NOTE: Make sure to completely review the Rubric for Milestone 2 Use the data from this Milestone and begin working on your final presentation due in Milestone 4 (Week 7) ITEMS TO COMPLETE FOR THIS MILESTONE (Green Tabs): GENERAL Use data from Milestone 1 in your analysis CONTRIBUTION MARGIN ANALYSIS Select a price for each service (grooming, day care, boarding) Determine the variable cost from the Variable_Fixed tab for each service Calculate the contribution margin for each service based on your sales price and the variable cost for that service BREAK-EVEN ANALYSIS Determine the fixed cost from the Variable_Fixed tab for each service Fixed & Variable cost designation is provided Calculate the break-even units (round up) for each service Calculate the break-even units (round up) for suggested target profit levels for each service
Paws and Company MILESTONE 1 - Variable & Fixed Cost Exercise Determine the per unit costfor each dog. Fill in the blanks to get the per unit cost and fixed cost of each service. Based on 5 grooms per day GROOMING Item Variable Costs Item Fixed Costs 2,080.01 Clipper(s) 1.37 30.95 0.72 Bowls 20.00 Towels 5.83 Utilities and Insurance 28.57 3.95 Scissors 1.01 Depreciation on heating system Cleaning Products: Odoban Cleaning Products: Simple Green 2.08 2.36 Draw 200.00 Total Variable Costs 9.97 Total Fixed Costs 2,367.92 k. Assume 25 operating days for kennel and 22 eight-hour work days for kennel attendant. DAY CARE Item Variable Costs Item Fixed Costs Day care attendant 2.88 232.14 Toys 0.27 150.00 0.10 Utilities and Insurance 214.29 Towels 1.40 Depreciation on heating system 29.64 Cleaning Products: Odoban Cleaning Products: Simple Green 15.59 16.78 Draw 200.00 4.65 Total Variable Costs Total Fixed Costs 858.44 Based on 12 dogs per day. Assume 22 eight-hour work days for kennel attendant. BOARDING Item Variable Costs Fixed Costs 5.62 Kennel attendant Depreciation on kennels 80.00 Bowls 0.86 Rent 386.90 Towels 0.87 250.00 Utilities and Insurance 357.14 Depreciation on heating system 49.40 Cleaning Products: Odoban Cleaning 25.98 Products: Simple Green 29.55 Draw 200.00 Total Variable Costs 7.35 Total Fixed Costs 1,378.97
Paws and Company MILESTONE 2 Contribution Margin per Unit & Contribution Margin Ratio INSTRUCTIONS: Select a price from the scenarios below and compute the contribution for each service based on your selected price. Variable cost per unit comes from your Variable_Fixed tab completed in Milestone 1. SCENARIO (choose one from each category): Dog Day Care 1. With pricing at $18 per dog per day, you can expect to have 22 dogs per day 2. With pricing at $20 per dog per day, you can expect to have 15 dogs per day 3. With pricing at $25 per dog per day, you can expect to have 10 dogs per day Overnight Boarding 1. With pricing at $25 per dog per day, you can expect to have 12 dogs per day 2. With pricing at $28 per dog per day, you can expect to have 10 dogs per day 3. With pricing at $30 per dog per day, you can expect to have 7 dogs per day Basic Groom 1. With pricing at $25 per dog per day, you can expect to have 5 dogs per day 2. With pricing at $30 per dog per day, you can expect to have 4 dogs per day 3. With pricing at $35 per dog per day, you can expect to have 3 dogs per day DAY CARE BOARDING GROOMING Sales Price per Unit Variable Cost per Unit Contribution Margin
Paws and Company MILESTONE 2 Break-Even Analysis INSTRUCTIONS: Show all steps and calculations to determine the break-even. Determine the break-even for the target profit levels as outlined in the instructions. Break-even Fixed Costs/ Contribution Margin DAY CARE BOARDING GROOMING Sales Price Fixed Costs Contribution Margin Break-even Units (round up) Target Profit 1,000.00 417.00 583.00 Break-even Units (round up) Target Profit 667.00 1,500.00 909.00 Break-even Units (round up)
Answer #1
Milestone 2
Day care Boarding Grooming
Sales price $18 $25 $25
Variable Cost 4.65 7.35 9.97
Contribution Margin 13.35 17.65 15.03
Milestone 2
Day care Boarding Grooming
Sales price 18 25 25
Fixed Costs 858.44 1378.97 2367.92
Contribution Margin 13.35 17.65 15.03
Break even units                       64                           78                      158
Break even units =
Fixed Cost / Contribution Margin
=858.44/13.35 =1378.97/17.365 =2367.92/15.03
Target Profit 417 583 1000
Break even units                       96                         111                      224
Break even units =
(Fixed Cost + Target profit ) / Contribution Margin
=858.44+417/13.35 =1378.97+583/17.365 =2367.92+1000/15.03
Target Profit 667 909 1500
Break even units                    114                         130                      257
Break even units =
(Fixed Cost + Target profit ) / Contribution Margin
=858.44+667 /13.35 =1378.97+909 /17.365 =2367.92+1500/15.03